Scenario 1:

You have to estimate the expected exchange rates one year from now between your home currency and the other currencies of the major other countries that you deal with in terms of both imports and exports. The reason is that increases in the values of other currencies compared to the U.S. Dollar may impact your imports negatively, whilst it may on the other hand, be good for exports. To do this estimate, you obtain the following spot exchange rate information:

£/$

0.76918

€/$

0.87616

You also obtain the following rates that you regard as similar to the annual risk free rates applying in the countries:

U.S.A.

2.660%

Britain

0.778%

France

0.500%

Your focus is presently to estimate the 12 month forward rates in order to consider the impact that it will have on the import and export sales of the company. Calculate the forward rates of the $ in terms of all the currencies by using simple interest rate parity e.g. 10% annual interest rate = 10/2 = 5% for six months. Do not apply effective annual interest rate compounding. Show all your workings in table 1 on the separate answer sheet by using the correct formula provided in your formula sheet.

Table 1: Calculation of 12 month forward rates using the simple interest rate parity principle (4 marks)

Exchange rate

Forward rate 12 months from now (provide answer in this column)

Workings (show calculations in this column)

£/$

€/$
Answer #1

£/$ forward rate = spot rate * ((1 + Britain interest rate) / (1 + US interest rate))

£/$ forward rate = 0.76918 * ((1 + 0.778%) / (1 + 2.660%)) =  £0.75508 / $

€/$ forward rate = spot rate * ((1 + France interest rate) / (1 + US interest rate))

€/$ forward rate = 0.87616 * ((1 + 0.500%) / (1 + 2.660%)) =  €0.85773 / $